Short Report – QA/QC Meeting September 2018

Meeting Date: 2018-09-20

Participants / Excused

LH, EPL, IB, TB, RM, MG, PM (7)  / IF, MS

Talking Points

Info

If you are (Tech-)SRP, please attend QA/QC meetings or tell LH if you can’t, needed for planning of the meetings.

SRP & Tech-SRP: please prepare short statement about your site and send to to LH (max. 2 sentences)

QA/QC Weekly checklist will be expanded: instead of adding Notes to the table, all technical Notes (errors) will be written directly in GIN. The link to the GIN entry will then be given in the table. This way all relevant information about sensor errors is collected in GIN.

  1. Do all SRPs have rights to add entries to GIN?
  2. Currently it is not possible to directly link to Fieldbook entries
  3. QA/QC issues need special tag in GIN
  4. new Logger (DAQ) checks: date, time check // webcam pics available // ping other devices in the network (only T-SRPs?) separate or current table?
  5. new Meteo checks: PWR voltage // internal battery voltage (check to avoid problems) // DNTP (date and time) separate or current table?

Reminders

  1. Automatic fluxes for last 14d (schedule is here)
  2. Online weekly QA/QC checks for nearly all sites
  3. Most recent eddy covariance raw data plots are found in the following folder on our group drive (P) (updated daily):

    P:\Flux\VPC_automatic_fluxes\vpc_automatic_flux_calcs/your_site/CO2*/EP-OUT*/plots/raw_data_csv_plots

Roadmap

  1. [IN PROGRESS] SRP QA/QC Weekly checks in online checklists (instead of the server Excel sheet) not implemented for CH-AWS
  2. [IN PROGRESS] T-SRP QA/QC Weekly checks in online checklists. T-SRPs will use the same list as SRPs, more checks will be added (?).
  3. [IN PROGRESS] Subscribe to your site RSS feed to get relevant news automatically. Please contact LH for subscribing to the feed so you can get the newest Swiss FluxNet news for your site. Active subscriptions at the moment: SRPs: CH-CHA (IF), CH-OE2 (RM), CH-DAV (MG), CH-FRU (LH); T-SRPs: CH-DAV (PM), CH-CHA (PL), CH-FRU (PL).
  4. [IN PROGRESS] On-demand flux calculations on grassland server dedicated ETH server instead of RDS. In progress. LH waiting for ISG.
  5. [UPCOMING] QV File Reports: checks available files on server. LH as soon as ICOS allows.
  6. [UPCOMING] GIN will be connected to SFN HP via RSS feeds. DO and LH will get in contact.

Site Reports

CH-AWS

SRP (LH, fluxes): Flux data are complete and look good.
Tech-SRP (IB): The 3D ultrasonic anemometer and the weather station are running normally. The adaptive mast is running, the snowheight sensor is almost ready to send snowheight data. The CO2/H2O sensor needs its thermocouple sensor replaced.

CH-CHA

SRP (IF): Reinstrumentation of the station (mainly soil plots) is planned soon, some of the meteo data shows bad values at the moment. Gaps in H2O and CO2 fluxes, probably because of dew formation on the open path sensor.
Tech-SRP (PL): PT100 of Netradiometer has sporadic appearing noise/false measurements and should be sent in for maintenance. PPFD_IN replica 1 of 2 is defective. Air pressure sensor has sporadic spikes. Markus is helping Iris to get Chamau up to date for her measuring campaign.

CH-DAV

SRP (MG, LH): Seems like the EC system stops logging after approx. 40 days, happened again on 8 Aug (WE implemented an additional script that should restart the EC system after such an error). Also, on of the soil sensors TS_1_6_1 seems to be damaged.
Tech-SRP (PM): Added atmospheric pressure sensor to embedded computer (additionally to datalogger). Data gap between August 8th and 9th.

CH-FRU

SRP (LH): Flux data look complete and good. Sensor G_GF1_0x06_2 has been sending erroneous data for some time and should probably be removed from the site.
Tech-SRP (PL): Power outage due to grid maintainance works on 26.09.2018. Our systems kept running thanks to the UPS. Else everything running smooth.

CH-INO NEW!

SRP (RM): Since the beginning of September sometimes more than four files are used for saving data on RAW. However, data looks complete.
T-SRP (MS): New compact meteo station for InnoFarm Project is installed is installed in OE for testing. Whole system runs on solar power. Everything works fine.

CH-LAE / CH-LAS

SRP (EPL): –
Tech-SRP (TB): (CH-LAE) TA/RH 27 and dendrometer needs data and on site checking. On the eddy data acquisition we have from time to time data leakage. -3.31 3.92 2.08 294.05 20.90 0.00 0.00 LI(1) 200 NO IRGA DATA! Problem is not easy to solve but is on the tasklist.

CH-OE2

SRP (RM): Between 7.9. and 20.9. two sensors (TS_CF4 and G_CF2) were not in the soil. Besides that, everything looks good.
Tech-SRP (MS): Heat Flux Plate and Soil Temperature) were dug out by someone. They were returned to their original position. New there is a WLAN access in OE2.

 

Snapshots

You are here: Short Report – QA/QC Meeting September 2018